### v1

# Data Caricatures: On the Representation of African American Language in Pretraining Corpora

Abstract
With a combination of quantitative experiments, human judgments, and qualitative analyses, we evaluate the quantity and quality of African American Language (AAL) representation in 12 predominantly English, open-source pretraining corpora. We specifically focus on the sources, variation, and naturalness of included AAL texts representing the AAL-speaking community. We find that AAL is underrepresented in all evaluated pretraining corpora compared to US demographics, constituting as little as 0.007% of documents. We also find that more than 25% of AAL texts in C4 may be inappropriate for LLMs to generate and reinforce harmful stereotypes. Finally, we find that most automated language, toxicity, and quality filters are more likely to conserve White Mainstream English (WME) texts over AAL in pretraining corpora.
[…]
.1em
Data Caricatures: On the Representation of African American Language in Pretraining Corpora
[…]
1 Introduction
Recent work in NLP has become increasingly interested in evaluating and mitigating African American Language (AAL) biases in generative language models Deas et al. (2023); Fleisig et al. (2024); Hofmann et al. (2024). AAL, the comprehensive and rule-governed language variety used by many, but not all, and not exclusively, members of the US African American community Grieser (2022), is known to be underrepresented in the C4 pretraining corpus Dodge et al. (2021). In particular, C4 is largely composed of White Mainstream English (WME, see Baker-Bell (2020)), the variety of English associated with White Americans and often found, for example, on Wikipedia. [^1] While other works in linguistics and NLP use ”African American Vernacular English (AAVE)” and ”Standard American English (SAE)” among other terms, we use AAL and WME to highlight the inherent interaction between race and language in social hierarchies. Additionally, we typically use ”language variety” to acknowledge the variation exhibited by AAL speakers across regions and contexts rather than a single, uniform dialect.
Figure 1: We evaluate the quantity and quality of automatically identified AAL documents in open source pretraining corpora. We find AAL is underrepresented across corpora and many documents contain texts that are misrepresentative of naturalistic AAL or reinforce stereotypes.
Given the close relationship between language and culture Adilazuarda et al. (2024), the distributions of sources that texts are drawn from may also misrepresent the broader culture of communities when used to train models. For example, many websites publish hip hop lyrics which may be collected and incorporated into pretraining data from online sources. While hip hop is an important component of Black culture and identity Clay (2003) as well as that of other communities, other registers of AAL (e.g., natural dialogue) are needed to more comprehensively capture the rich and diverse facets of Black culture. Furthermore, hip hop lyrics are not independently representative of AAL as a whole; lyrics do not necessarily reflect natural speech and no explicit indication of genre or register is provided during pretraining. Prior studies validate this gap, finding that Black users perceive a lack of cultural knowledge in current language technologies; they often suspect that AAL speakers are overlooked in model development processes Brewer et al. (2023); Cunningham et al. (2024).
Therefore, in addition to the quantity of AAL that is represented, we argue that the quality of the representation of AAL is also important. [^2] Note, biases may also come from the technical representation of text in the form of text embeddings. This is distinct, however, from the representation that we explore in this work, which refers to how the language of AAL speakers is depicted in pretraining corpora. We define quality as the extent to which texts faithfully and authentically portray the linguistic patterns–and consequently culture–of native AAL speakers. Accordingly, low quality representations of AAL (e.g., texts written by non-native speakers) pose representational harms to AAL speakers Blodgett et al. (2020).
We focus on three guiding research questions:
  1. RQ1)
    What percent of documents in modern, open-source pretraining corpora contain AAL and specific AAL features?
  2. RQ2)
    What is the quality of the included AAL texts in terms of diversity, authenticity, and lack of offensiveness?
  3. RQ3)
    How do modern data quality filtering strategies behave with AAL texts?
We use mixed-methods studies to address these questions, incorporating quantitative experiments, human judgments, and qualitative examination of the documents in pretraining corpora. To investigate RQ1 (Section 4), we measure the prevalence of AAL in 12 open-source pretraining corpora and find that as few as 0.007% of pretraining corpora documents contain AAL. To investigate RQ2 (Section 5), we conduct automated and human evaluations of AAL texts in pretraining corpora and find that a significant number of documents are unrepresentative of naturalistic AAL [^3] By naturalistic AAL, we refer to the linguistic patterns in text that most closely resemble everyday, spontaneous speech. (e.g., perceived to be written by non-native speakers). Surprisingly, some texts resembling AAL are actually posted by corporate social media accounts. Finally, to investigate RQ3 (Section 6), we evaluate 16 quality filtering approaches and find that most filters are more likely to conserve WME texts compared to AAL, particularly for texts from social media or song lyrics.
2 Background and Related Work
AAL, Performativity, and Misrepresentation. Research involving AAL in the NLP community has relied on texts drawn from different sources, such as speech transcripts Farrington and Kendall (2021) and social media Blodgett et al. (2016). Across such sources, AAL use can reflect varying degrees of what has been termed performative speech, language practices which use the rhetorical style of AAL such as signifyin’ Mitchell-Kernan and Thomas (1972) on a speaker’s cultural background to construct or communicate Black identity. Because they primarily rely on language to communicate Black culture, performative register is especially prevalent in hip hop Alim (2006) or social media language Eisenstein (2013); Ilbury (2020). Recent work has also argued that the linguistic patterns of these sources may misrepresent the patterns of AAL-speakers’ when interacting with LLMs Kleiner et al. (2024). Alternatively, AAL can be misrepresented through the derogatory use of features Ronkin and Karn (1999), diffusion of linguistic markers Corradini (2024), and use of AAL-like language in marketing Roth-Gordon et al. (2020). In this work, we examine the quality of AAL representation by considering performativity and these potential misrepresentations.
AAL Biases in Language Technologies. The most prominent documentation and mitigation of AAL biases in text-based language models has focused on toxicity detection (e.g., Sap et al. 2019a; Harris et al. 2022; Davidson et al. 2019; Cheng et al. 2022; Halevy et al. 2021). More recent work evaluates AAL biases in other contexts, including classification tasks through synthetic data Ziems et al. (2022); Dacon et al. (2022), extractive summarization Keswani and Celis (2021); Olabisi et al. (2022), LLMs’ stereotyping behaviors Fleisig et al. (2024); Hofmann et al. (2024), and generative language models Groenwold et al. (2020); Deas et al. (2023, 2024). Rather than evaluate downstream biases, we specifically examine AAL in pretraining data as a potential source of bias.
Pretraining Data Quality. Hovy and Prabhumoye (2021) identify data as a fundamental source of bias, and prior work has attempted to trace various biases to training data (e.g., Ladhak et al. (2023); Feng et al. (2023); Dodge et al. (2021)). Recent work has also increasingly been concerned with measuring and selecting high quality data for pretraining LLMs (e.g., Wettig et al. (2024)). Such studies, however, use a variety of different measures of data quality. A majority of prior work (e.g., Li et al. 2024) primarily operationalizes quality through downstream model performance on broad benchmarks (e.g., MMLU Hendrycks et al. (2021) or HellaSwag Zellers et al. (2019)). Other work measures quality using sources that are typically considered "high-quality" (e.g., Wikipedia) in developing automated data filters (e.g., RedPajama, Computer (2023); The Pile, Gao et al. (2020)). Finally, relatively fewer studies have examined intrinsic notions of quality, such as educational value or lack of harmful content Wettig et al. (2024); Sachdeva et al. (2024); Gunasekar et al. (2023). With the aim of identifying potential causes of AAL biases, we measure quality through a variety of means, focusing on texts’ sources, representativeness of naturalistic AAL, and lack of harmful reflections on AAL-speakers.
3 Data Extraction
Dataset Example Models # Docs (sample) % CC Docs # AAL Docs (%)
Full Corpora
OpenWebText Gokaslan et al. (2019) SSDLM, RoBERTa 8M 0% 2,684 (.03%)
The Pile Gao et al. (2020) GPT-Neo, GPT-J 140M 3% 109,333 (.08%)
Dolmino (Dolmino-mix) OLMo et al. (2024) OLMo-v2 165M 83% 53,937 (.03%)
C4 Raffel et al. (2020) T5 365M 100% 280,604 (.07%)
C4.NoBlockList Dodge et al. (2021) N/A 395M 100% 447032 (.11%)
RefinedWeb Penedo et al. (2023) FalconLM 968M 100% 1,143,244 (.12%)
RedPajama Computer (2023) OpenLlama 968M 88% 63,189 (.007%)
FineWeb-Edu Penedo et al. (2024) N/A 1.8B 100% 15,907 (.0009%)
Dolma Soldaini et al. (2024) OLMo-v1 2.5B 78% 3,119,429 (.12%)
Sampled Corpora
Dolmino (Olmo-mix) OLMo et al. (2024) OLMo-v2 3.1B (238M) 96% 317,178 (.02%$\pm$.001%)
DCLM-Baseline Li et al. (2024) DCLM-Baseline 3.2B (105M) 100% 11,673 (.01%$\pm$.004%)
RedPajama-v2 Weber et al. (2024) N/A 20.8B (176M) 100% 684,099 (.18%$\pm$0.09)
FineWeb Penedo et al. (2024) N/A 48.6B (38M) 100% 10,470 (.03%$\pm$.005%)
Table 1: The 12 open-source pretraining datasets evaluated, including example models pretrained on each dataset, size in raw unique documents, and proportion composed of Common Crawl. indicates only non-copyrighted portions are available and included in analyses. Bolded corpora indicate that C4 is explicitly included in the corpus before further filtering. Given their size, the bottom 4 corpora are analyzed using a random sample of dataset shards; 99% confidence intervals on each sampled corpora’s estimate is included.
We examine the quantity and quality of AAL texts present in 12 predominantly English, open-source pretraining corpora. All corpora are listed in Table 1. In particular, we consider FineWeb-Edu (italicized in Table 1) as a baseline, given that it strictly prioritizes highly educational content (e.g., Wikipedia) as opposed to diverse everyday language use. Following prior work’s documentation of C4 Dodge et al. (2021) and because many corpora predominantly rely on Common Crawl texts, we focus our human judgments and AAL feature analyses on variants of the C4 corpus Raffel et al. (2020).
3.1 Extracting AAL Subsets
To identify AAL texts, we follow Dodge et al.’s (2021) analysis of C4 as well as other related work (e.g., Xia et al. (2020); Sap et al. (2019b); Davidson et al. (2019)) and use a mixed-membership demographic-alignment classifier validated with common linguistic features of AAL Blodgett et al. (2016). From eight full corpora, we extract all texts where AAL is the most likely classification (additional corpora details are included in Subsection A.1). For four exceedingly large corpora with more than 3 billion documents–Dolmino/Olmo-mix OLMo et al. (2024), DCLM-baseline Li et al. (2024), FineWeb Penedo et al. (2024), and RedPajama-v2 Weber et al. (2024)–we analyze a 250 GB sample of each corpus. Additional extraction methodology and corpus details are included in Subsection A.2 and Appendix B.  [^4] While corpora like RedPajama-2 are intended to enable experimentation with automatic filtering rather than to be used as is for pretraining, we refer to all corpora as ”pretraining corpora” for simplicity. To account for AAL features that may be present within documents that largely reflect WME, we additionally extract a more conservative subset of AAL texts using a threshold of 0.3 for these analyses.  [^5] While prior work largely considers a threshold of 0.8 (e.g., Xia et al. (2020)), we choose a threshold of 0.3 to calculate a more conservative estimate of feature prevalence while maintaining a manageable corpus size for analysis. Additional discussion is included in Appendix C.
We use the same block list of terms used in the original C4 corpus Raffel et al. (2020) to identify texts that would be filtered from the corpus before pretraining. This filtering yields two distinct variants: C4.en.noBlocklist which does not employ the block list, and C4.en which lacks block list-identified texts. With these two variants, we conduct a fine-grained evaluation of the block list’s impacts on AAL representation, prevalence of features, and human judgments of AAL texts.
4 RQ1: How Much is AAL Represented?
4.1 Methods
Human Judgments. To complement automatic analyses, we collect human judgments of randomly sampled C4 texts within the subset of automatically extracted AAL texts. In sampling texts for judgments, we prioritize texts with higher probability of containing AAL according to the demographic alignment classifier as well as texts from C4.en.
We recruit three annotators to conduct the human judgments of the sampled texts; all annotators are self-reported native AAL-speakers. Annotators also currently study or have previously studied linguistics or computational linguistics. Annotators are provided a text from either the AAL subset of C4.en.noBlocklist or C4.en and asked to label each text first on two dimensions drawn from prior work Deas et al. (2023): Human-Likeness asks whether the text appears authored by a human and Linguistic Match asks whether there are identifiable features of AAL in the given text. Both dimensions are judged on 4-point Likert scales (higher representing more human like and more identifiable features). Among all three annotators, we collect judgments for 1,054 texts. Annotators exhibit moderate to substantial agreement for binarized Human-Likeness and Linguistic Match dimensions ($\kappa=0.581$ and $\kappa=0.747$ respectively). Additional sampling, annotation, and interface details are included in Appendix D.
Automated Feature Extraction. To conduct a more fine-grained analysis, we estimate the distribution of morphosyntactic AAL features. Features are identified using the CGEdit model Masis et al. (2022), a classifier developed with a human-in-the-loop framework. The classifier considers 17 AAL morphosyntactic constructions, including features like habitual be (e.g., "He be driving") and copula deletion (e.g., "He $\emptyset$ at home"). See Masis et al. 2022 and Appendix E for a detailed list of features. As the classifier was shown to be reliable for examining distributions in large collections of text, we restrict analysis of features to trends in large data subsets rather than individual texts.
4.2 Results
AAL Frequency. Table 1 includes the proportion of documents labeled as AAL in each corpus evaluated. Language understanding benchmarks drive pretraining data curation choices and are drawn from sources such as descriptive video captions (e.g., HellaSwag Zellers et al. (2019), Swag Zellers et al. (2018)) or academic exams (e.g., MMLU Hendrycks et al. (2021)). Therefore, we expect that the analyzed corpora contain few documents with AAL. In line with this intuition, we find that AAL is extremely underrepresented across corpora used and intended for LLM pretraining, included in as few .007% of documents (i.e. RedPajama). As expected, the FineWeb-Edu corpus, which employs among the strictest filtering approaches, exhibits the lowest percentage of documents containing AAL (.0009%), while RedPajama-v2, which employs relatively limited filtering, exhibits the highest percentage (.18%). Due to this trend and those found in prior work, we investigate filtering approaches further in Section 6.
AAL Prob. C4 # Docs % AAL Judgments
$0.5\leq p\leq 0.6$ 41,930 44.7%
$0.6\leq p\leq 0.7$ 12,913 36.3%
$0.7\leq p\leq 0.8$ 4,319 36.7%
$0.8\leq p\leq 0.9$ 922 30.9%
$0.9\leq p$ 120 23.0%
Table 2: Percent of texts labeled by human annotators as containing AAL features for each range of posterior probabilities of the demographic alignment classifier.
To verify the automated analyses, Table 2 presents the proportion of texts labeled by annotators as containing identifiable AAL features within different ranges of AAL posterior probability as output by the demographic alignment classifier. Annotators indicate that a significant proportion of the texts do not contain features of AAL (55.3% in the $0.5\leq p\leq 0.6$ range). Many of these documents instead contain lists of search terms and artist names that resemble lexical markers of AAL (e.g., Lil Wayne) rather than naturalistic AAL.
AAL Feature Frequency. Figure 2 presents the frequency of each automatically detected AAL feature in C4.en.noBlocklist and C4.en. Supporting Dodge et al.’s (2021) finding that the block list disproportionately filters AAL texts, we see that filtering lowers the frequency of all AAL features. Some of the most frequent features in C4.en.noBlocklist–zero copula (ZC) and multiple negatives (MN)–are among the most frequent documented features of naturalistic AAL. Other features such as negative auxiliary inversion (NAI) and zero plurals (ZP), however, have extremely low representation despite being relatively common features among both urban and rural AAL-speakers Kortmann et al. (2020).
While some features are represented in frequencies that generally reflect naturalistic AAL, Figure 2 also illustrates how these frequencies are impacted by filtering. While the most frequent features are generally shared among both subsets, the order of most frequent features diverges. While Zero Copula (ZC; e.g., he $\emptyset$ running) is detected most often in C4.en.noBlocklist, it is far less represented in C4.en. This alteration suggests that filtering may impact AAL-speakers differently depending on region, class, and other speaker characteristics (e.g., Southern or working class; Wolfram and Kohn 2015). Not only does C4 filtering disproportionately remove AAL texts, but filtering also systematically alters the distribution of AAL features.
Figure 2: Document-level frequency of morphosyntactic features among texts in C4.en and C4.en.noBlocklist. Features ordered by frequency in C4.en. List of feature abbreviations and descriptions are included in Appendix E.
5 RQ2: How Well is AAL Represented?
5.1 Methods
Human Judgments. For each text that is judged to be human written and contain identifiable AAL features, annotators are asked to also rate each text on three additional dimensions. First, we follow Deas et al. (2024) and include Native Speaker which asks whether the text appears to be naturally written by an AAL-speaker. The remaining two dimensions capture whether the AAL texts are suitable for training LLMs. Inspired by Fleisig et al. (2024), Stereotype asks whether the text portrays a stereotypical or harmful representation of AAL or its speakers. Finally, we consider annotators’ own linguistic attitudes toward texts that are included in pretraining corpora, motivated by sociolinguistics work on listeners’ language attitudes and perceptions Campbell-Kibler (2008). We include Appropriateness which asks whether an annotator perceives a given text to be appropriate for language models to generate and thus appropriate to include in pretraining data.
All dimensions are again judged on 4-point Likert scales with higher values representing that a text is judged to be likely written by a native speaker, to perpetuate stereotypes, and to be appropriate for an LLM to generate. Annotators exhibit substantial agreement for Native Speaker ($\kappa$ = 0.619) although annotators’ varying personal perspectives on the Appropriateness and Stereotype dimensions expectedly yield little to no agreement ($\kappa=$ 0.188 and -0.021 respectively), highlighting the need for representation of AAL-speakers in model development processes. Additional details and discussion of agreement are included in Appendix D.
Misrepresentative Language. To explore whether the representation of AAL in C4 is indicative of naturalistic AAL, we first estimate the presence of hip hop and rap lyrics in C4 variants. We particularly focus on song lyrics because large-scale corpora of published hip hop lyrics are available and enable measuring overlap with each pretraining corpus. Leveraging the aforementioned human judgments, we also estimate the perceived presence of AAL use by non-AAL-speakers.
We follow the deduplication approach of Brown et al. 2020 which measures the overlap of 8-13 token n-grams. While an ideal analysis would measure exact matches, through manual examination, we find instances of hip hop lyrics embedded in non-lyric text, modifications to lyrics such as censored terms, and orthographic differences (e.g., "going" vs. "goin") that would not be captured by such an approach. We note that while we use a large corpus of song lyrics, a comprehensive dataset remains infeasible and our estimates likely reflect lower bounds. Methodological details are included in Appendix F and additional analyses are included in Appendix G.
5.2 Results
Figure 3: % of AAL documents extracted from C4 containing shared n-grams with hip hop lyrics for varying n-gram token lengths. Tokenization uses the Llama-2 tokenizer.
C4 Subset Text
EN You go dey hear: ha ha! Catch am, catch am! Thief, thief, thief! Catch am, catch am! Rogue, rogue, rogue!
Cant Say My Name But Rap About A N*ggas Wife. You So Black & White Tryna Live A N*ggas Life… You Aint Wettin Nobody You Canady Dry
EN.NoBlockList Thank Racks on racks on racks on racks / None of my cars ain’t rented, all mine black, my windows tinted
Table 3: Examples of hip hop lyrics found in C4.en and C4.en.noBlocklist. Texts are shown exactly as they appear in the corpus.
Hip Hop Lyrics. Figure 3 presents the percentage of documents overlapping with hip hop lyrics by n-gram token length. Nearly 15% of C4.en.noBlocklist and 12% of C4.en documents contain substantial overlap (8-grams) with hip hop lyrics. Examples of lyrics identified by annotators are included in Table 3. While there are AAL features present, such as the use of negative concord in None of my cars ain’t rented, any inclusion of hip hop lyrics may be misrepresentative of how AAL-speakers would interact with LLMs.
Non-AAL-Speakers. AAL may also be misrepresented through use by non-AAL-speakers, predominantly online. While it is infeasible to automatically determine the linguistic background of a text’s author, among the C4.en texts annotators judged to contain AAL features, 51% were judged to have been unlikely written by a native AAL-speaker. Most of these texts (70%) were perceived as not human-written.  [^6] As an approximate reference, only 21% of human-written AAL texts in Deas et al. (2023) were judged to be likely not human-written or unnatural. The remaining texts, however, include corporate social media accounts adapting AAL-like language (e.g., …this will get you where you need to be. Ball out by clicking the link below.) as well as online forum posts (e.g. Dat be that real deal Hip Hop! …Put some spect on it!). These texts often exaggerate the use of AAL features, contributing to misunderstandings of native speakers. We conclude that a significant portion of texts with AAL features are unlikely to have been written by an individual AAL-speaker, and misrepresent naturalistic AAL.
Many features of AAL are shared with other language varieties and the appearance of shared features in different linguistic contexts may complicate LLMs’ ability to model these varieties during pretraining. Annotators noted many of these cases, such as the Hawai’ian Creole text …An make us shame cuz we no mo husban. Dat Yahweh make come up from da king ohana… which may be misinterpreted as Digital AAL (DAAL; Cunningham (2014)). [^7] See Appendix H for further discussion.
Figure 4: Proportion of extracted AAL from C4 judged to be inappropriate for models to generate and judged to reinforce stereotypes of AAL or its speakers. Proportions are show separately for the filtered and unfiltered subsets of the extracted data.
[…]
…Been ****in wit jack and the mob for the longest. been ****in wit this MOB muzic 4 a very long time , Jack!…" 0 2
He gon fuk around & drown off this ? 0 0
Drake is two levels up and Meek Mill is still not off the mark. The Toronto rapper just dropped another diss song “Back To Back” where he freestyle about his former friend and now nemesis. 3 1
Table 4: Examples of AAL texts from C4.en with appropriateness (A) and stereotype (S) judgments.
Innapropriateness and Stereotyping. The results of Appropriateness and Stereotype judgments are shown in Figure 4. First, we find that the filtering applied to C4.en.noBlocklist removes many inappropriate and stereotype-reinforcing texts as intended, with C4.en.noBlocklist exceeding the rates of C4.en. We also find, however, that substantial proportions of remaining texts in C4.en are judged to be inappropriate and stereotype-reinforcing–28.9% and 24.5% respectively. Many of these texts contain variants of C4 block list terms or terms self-censored with asterisks. For example, the first two texts in Table 4 are judged to be inappropriate for models to generate and contain references to block list terms (****in and fuk).
6 RQ3: How does filtering impact AAL representation?
The previous experiments show that AAL is often underrepresented and misrepresented in modern, open-source pretraining corpora. We follow these experiments with an evaluation of automated filtering strategies used in pretraining corpora to better understand potential underlying determinants of AAL representation.
6.1 Methods
Source Dialect # Documents Avg. Length
RedPajama-v2 AAL 235,490 509.8
WME 784.7
  Dialogues AAL 11,787 59.9
Song Lyrics AAL 10,000 441.0
Social Media AAL 50,000 19.7
Table 5: Summary of datasets used in filtering experiments. We sample equal numbers of texts to represent AAL and WME, and where possible, sample texts of similar length.
The C4 block list is known to disproportionately filter AAL from pretraining data Dodge et al. (2021). To investigate how recent, model-based language, toxicity, and quality filters affect AAL representation, we evaluate 16 automated filters on AAL texts. A full list with descriptions of each filter evaluated is included in Appendix I.
We conduct two sets of experiments to evaluate the behavior of automated filters on AAL texts. First, we evaluate automated filters on AAL and WME texts in RedPajama-v2 Weber et al. (2024) to examine how filters treat web texts that naturally occur in sources of pretraining data. While this experiment reflects real use cases of filters, it does not provide insights into behavior across varying sources of AAL texts. In a second controlled experiment, we collect AAL from different sources (dialogues, hip hop lyrics, social media) to evaluate how filters’ behaviors differ among domains.
RedPajama Filtering RedPajama-v2 Weber et al. (2024) is a large-scale pool of Common Crawl text with pre-computed filtering heuristics. It was introduced to enable efficient filtering rather than to directly pretrain LLMs without further pre-processing. To evaluate automated filters, we extract all texts with an African American-aligned English probability $\geq 0.8$ and randomly sample an equal number of texts with a White-aligned English probability of $\geq 0.8$ following prior work Xia et al. (2020).
AAL Source Filtering. To analyze different sources of AAL, we collect AAL texts from different domains (as in Deas et al. (2023)). We evaluate automated filtering on three different domains of AAL-like language: social media, hip hop lyrics, and dialogues. For social media texts, we use the African American-aligned subset of the TwitterAAE corpus Blodgett et al. (2016). For hip hop lyrics we randomly sample full song lyrics from hip hop and rap songs. Finally, for dialogues, we use CORAAL Farrington and Kendall (2021) to represent naturalistic AAL.  [^8] Consecutive turns by the same speaker are merged to account for dialogue turns consisting of few tokens (e.g., ”Yes”). A summary of all filter evaluation datasets can be found in Table 5.
Filter Evaluation. We run each filter on the aforementioned datasets for evaluation. We compare the z-score normalized ($z=\frac{x-mean}{stdev}$) predictions for each automated filter based on its form: probabilities for classifiers, perplexities for n-gram models, and ratings for LLM-as-a-judge approaches. For all classifiers and rating-based models, we consider the score given to the positive class or direction such that higher scores represent a higher likelihood of text being preserved. We negate perplexities for applicable filters such that higher z-scores similarly indicate a higher likelihood of being conserved. We assess significant differences using two-tailed t-tests of the means.
6.2 Results
Figure 5: Average normed filter probabilities given by each automated filter for the positive label across AAL and WME in RedPajama (top) and across data sources (bottom). Error bars represent 99% confidence intervals, and (top) indicates significant differences ($p<0.01$). Shaded columns indicate filters that prefer AAL texts (top) and prefer Hip Hop/AAL Tweets (bottom). Larger values indicate texts are more likely to be preserved.
RedPajama-v2. We first examine how automated filters behave with AAL and WME data distributions similar to those in pretraining corpora using RedPajama-v2 Weber et al. (2024). Figure 5 (top) presents the normed average filter outputs for the positive label (i.e., the label more likely to conserve a given text) on the AAL and WME subsets of RedPajama-v2. Most filters (13 of 16 ) are more likely to remove AAL texts than WME texts. Supporting prior work, language Blodgett and O’Connor (2017) and toxicity Sap et al. (2019b) filters are included among those that assign lower scores to AAL texts. Interestingly, two of the remaining filters (Wiki and Wiki vs. RW) include Wikipedia data in the "high-quality" reference texts used to develop the filter.
AAL Sources. To better understand how automated filters’ predictions vary across AAL sources, Figure 5 (bottom) presents the same average normalized model predictions on AAL texts from dialogues (CORAAL), song lyrics (Hip Hop), and social media (AAL Tweets). As CORAAL contains natural speech and is unlikely to contain the lexical features exhibited by hip hop lyrics and social media, most filters (11 of 16 ) expectedly are more likely to conserve dialogue transcripts over the other sources. Most filters (12 of 16 ) are also more likely to conserve AAL tweets over hip hop lyrics. While filters appear to elevate natural speech, texts like transcripts in CORAAL are not widely available online suggesting that the sources of pretraining data may be the more influential factor determining AAL’s representation in corpora (see Appendix J for further discussion).
7 Discussion and Conclusion
In this work, we examine the quantity and quality of AAL representation in pretraining corpora as well as the impact of recent automated quality filters. We show that AAL is widely underrepresented in modern corpora, and its representation is highly sensitive to choices made during dataset construction–namely data sources and filtering. Underlying these findings, language understanding benchmarks that drive LLM development are unlikely to contain AAL, potentially incentivizing the similarly minimal AAL representation found in pretraining corpora. Poor representation may reinforce the discriminatory language model behaviors found in prior work Fleisig et al. (2024); Hofmann et al. (2024) and restrict the benefits of LLMs for AAL-speakers Brewer et al. (2023).
Driven by our findings and prior work, incorporating understanding of text sources in LLM development and carefully curating these sources through inclusion of AAL-speakers as stakeholders in design processes Friedman (1996); Suresh et al. (2024) is necessary to faithfully represent naturalistic AAL use. We recommend that in developing LLMs designed for socially impactful domains where AAL-speakers already face disparate impacts (e.g., medical Beach et al. (2021) or legal Jones et al. (2019) settings), pretraining data curators ensure that the sources and filtering processes retain representative AAL data. Such precautions may also accordingly be extended to the benchmarks used to evaluate the effectiveness of design decisions throughout the LLM development cycle, incentivizing understanding of sociolinguistic variation as well as commonsense understanding and other capabilities.
8 Limitations
We outline limitations to consider alongside our presented results. First, there are a wide variety of pretraining corpora and filtering strategies in use for LLM development, and the corpora we evaluate are not exhaustive. Notably, our work is enabled by the efforts to open-source language model development, but many corpora, such as those used to pretrain GPT-4o OpenAI et al. (2024), Llama-3 Grattafiori et al. (2024), and similarly large, closed-source models are unavailable for analysis. We include multiple corpora in our experiments, particularly those that attempt to replicate commercial corpora to better ensure that the findings are broadly applicable to current and future models. Additionally, given that online texts are the primary source of data for pertaining, we focus experiments on corpora composed of Common Crawl and internet sources.
Furthermore, we collect human judgments of AAL text quality, though we acknowledge that our recruited annotators may not be fully representative of the perspectives and opinions of the broader AAL-speaking community. We recruit annotators that have experience in linguistics, NLP, and AAL as they are equipped to identify AAL features and consider the implications of using specific texts in LLM development. We hope that future work will consult diverse and representative annotators in both analyzing and curating data to better inform pretraining data collection.
Finally, we follow prior work in using the demographic alignment classifier from Blodgett et al. (2016) to extract AAL from pretraining data for analyses. The classifier, however, is fitted on Twitter data which may not generalize to other language contexts. We broaden our analysis to consider any texts with higher than 0.3 probability of containing AAL to mitigate the impacts of the potential domain difference. Similarly, the CGEdit classifier was not initially developed for internet texts. However, analyses using CGEdit focus on morphosyntactic constructions which are similarly portrayed online and in speech unlike features linked to orthography (e.g., Eisenstein 2013).
9 Ethics Statement
Author Positionality.
As a significant portion of this work involves qualitative analyses of texts, estimates, we acknowledge that the authors’ backgrounds and perspectives may impact interpretation of some findings. In this work, we attempt to calculate and report quantitative statistics to complement qualitative discussion of pretraining corpora’s contents. As we strictly analyze open-sourced pretraining corpora, we encourage readers to examine texts within each corpus as well as form their own interpretations from the presented results.
AAL in Pretraining and Benchmark Data.
Our analyses show that AAL is underrepresented in all pretraining corpora studied, and that many documents are misrepresentative of naturalistic AAL. While increasing the representation of AAL in these corpora may improve downstream understanding in LLMs, at the same time, increased representation can pose risks to AAL speakers. As discussed by Patton et al. (2020), consideration of the context of data and impact of technologies on vulnerable populations is necessary for research involving these communities. Particularly if data is collected without the consent and inclusion of AAL-speaking communities, increased representation in corpora may increase privacy risks Brown et al. (2022) or perpetuate misrepresentative notions of AAL by means of LLMs.
AAL-use by LLMs.
Our findings contribute to the ongoing study of AAL biases in both language understanding and generation systems. Prior work highlights the barriers that language technologies’ lack of AAL understanding poses to African American users Cunningham et al. (2024), while others detail how human prejudices can lead to, for example, disproportionately labeling AAL as toxic Sap et al. (2022). With respect to language generation, a recent study has found that African American users prefer chatbots that use WME rather than AAL in generations Finch et al. (2025). These findings highlight the conflict between language technologies that are able to properly understand AAL but not necessarily to generate AAL. As such, while our findings document factors in pretraining data that may lead to misunderstanding of AAL in LLMs, increasing representation alone does not absolve language model creators of their obligation to take in the culturally relevant context.
To conclude, no LLM depiction of AAL should misrepresent and caricature AAL use. In contexts where LLM use serves the social benefit of AAL-speakers, measures should be taken to ensure that depictions of AAL do not trivialize the rich linguistic and cultural background surrounding them.
Appendix A Large-Scale Corpora Analysis
We select 12 open-source pretraining corpora for our analyses, prioritizing those that have been used to pretrain models often used in academic research. Given our focus on AAL, we also focus analyses corpora that are intended for pretraining predominantly English rather than multilingual models. For all corpora, we strictly analyze portions that are publicly available and do not include subsets of RedPajama Computer (2023) or The Pile Gao et al. (2020) that are restricted due to copyright. All corpora are only used for academic research purposes in line with their intent.
A.1 Pretraining Corpus Scanning
Our data pipeline is written in Python, making use of the Polars library Vink et al. (2025) for performance-sensitive data processing, as well as Pandas pandas development team (2020) and MatPlotlib Hunter (2007) for analyses and visualization. Labeling of independent subsets of the data is a trivially parallelizable task. We partition the corpus into disjoint splits and use 20 independent worker processes, one for each available physical CPU core to store the full text of records matching our selection criterion. In all, we obtain labels for over 16 TB of compressed records, as stored on disk. For capturing approximations of full distributions, we additionally use $t$-digest Davidson-Pilon (2015).
A.2 Corpus Sample Analyses
Our file sampling policy was motivated by the constraints of the available metadata, the size of the corpora, and the goal of our analysis being reproducible. File sizes varied between the corpora and within corpora, so a fixed number of files would create an in. With the sheer amount of variance in the metadata of the corpora, the only piece of metadata common to all corpora was the URL hosting the data file. To create a uniform random sampling of the files in the corpora, we use a variant on hash-based consistent sampling Manasse et al. (2010). We hash the full url hosting the data with sha256, sample the first $n$ files sorted by hash greater than our threshold of 250 GB per corpus, due to constraints on our resources.
Appendix B Additional Data Details
Figure 6: Distribution of metadata from selected corpora.
Appendix C AAL Probability Threshold
[…]
Figure 7: Full distribution of demographic alignment classifier posterior probabilities across all TwitterAAE Model Labels for C4 and all sampled corpora datacomp-lm-baseline, Redpajama-Data-V2, and fineweb, as captured by t-digest Davidson-Pilon (2015).
Any record where AAL has the greatest probability of all labels or where AAL exceeds a posterior probability of 0.3 is considered for analysis. This threshold is chosen for the individual feature analyses to capture documents that may contain features of AAL although largely be composed of WME. At the same time, this threshold does not add significant documents to the extracted dataset that would make analysis infeasible with the available computational resources. Figure 7 shows the distribution of document AAL probabilities as well as the probabilities of the other language varieties considered by the demographic alignment classifier Blodgett et al. (2016). For C4 (top), considering lower thresholds sharply increases the number of documents considered.
Appendix D Human Judgments
D.1 Annotation Details
We recruit 3 annotators that self-identify as AAL-speakers and have previous or current experience in natural language processing or linguistics. Each annotator is initially provided 420 texts total, with 50 texts shared among all annotators. Two annotators completed judgments for all 420 texts, and the remaining annotator completed 299 texts. Annotators were compensated at a rate of \$22.50 per hour.
D.2 Human Judgments Sample
Because much of recent work relies on a threshold of 0.8 to identify AAL documents (e.g., Davidson et al. (2019)) using the demographic alignment classifier Blodgett et al. (2016), we prioritize texts with a higher probability of including AAL in human judgments. Additionally, because C4.en.noBlocklist is not used for pretraining, we prioritize texts from the filtered subset of C4 (C4.en). The composition enforced on the sampled human judgment texts is shown in Table 6. Notably, only 10% of the texts in the sample have $\geq 90\%$ probability because there are few texts in the full extracted dataset that meet this requirement. Considering this sampling, all reported estimates based on human judgments (e.g., overall Stereotype estimates) are calculated through a weighted average across probability ranges, weighted by the size of each probability range in the full corpus.
AA Prob % Sample
.$.9\leq P_{\_}{AA}.$ 0.1
.$9\displaystyle.8\geq P_{\_}{AA}\geq.$9 0.35
.$8\displaystyle.7\geq P_{\_}{AA}\geq.$8 0.3
.$7\displaystyle.6\geq P_{\_}{AA}\geq.$7 0.15
.$6\displaystyle.5\geq P_{\_}{AA}\geq.$6 0.1
Filtering % Sample
Filtered (C4.en) 0.75
Unfiltered (C4.en.noBlocklist) 0.25
Table 6: Percent sampled from each subset of the extracted AAL C4 data for human judgments.
D.3 Annotator Agreement
Dimension $\kappa$ $\kappa_{\_}{bin}$ $r_{\_}s$ Support
Human 0.226 0.581 0.588 40.3
Dialect 0.357 0.747 0.699
Native 0.095 0.619 0.572
Approp 0.101 0.188 0.336 13.7
Stereo 0.009 -0.021 0.083
Table 7: Average pairwise interannotator agreement results for each dimension using Cohen’s $\kappa$ and Spearman $r_{\_}s$. Cohen’s $\kappa_{\_}{bin}$ represents agreement with binarized judgments ($\{0,1\}\rightarrow 0$, $\{2,3\}\rightarrow 1$).
Table 7 presents the inter-annotator agreement scores for each judgment dimension. Annotators show moderate to substantial agreement for binarized Humanness ($\kappa_{\_}{bin}=0.581$), Dialect ($\kappa_{\_}{bin}=0.747$), and Native Speaker ($\kappa_{\_}{bin}=0.619$), as well as strong Spearman correlations.
Agreements on the two remaining perceptual dimensions, Appropriateness and Stereotype, are expectedly low considering that annotators personal linguistic attitudes can lead to varying interpretations of what constitutes "appropriate" or appears to reinforce stereotypes. However, annotators exhibit slight agreement on Appropriateness.
D.4 Annotation Interface
Because we are interested in capturing the annotators’ own perspectives on the pretraining texts and their interpretations of these dimensions, we avoid providing detailed definitions of the dimensions–particularly the Stereotype and Appropriateness dimensions. We also include an additional space for comments, including reasoning for selected judgments as well as if the annotator recognizes a given text from a different source. Screenshots of the interface are provided in Figure 8, Figure 9, and Figure 10.
Figure 8: Screenshot of annotation interface instructions panel.
Figure 9: Screenshot of Human-Likeness, Linguistic Match, and Native Speaker questions in annotation interface.
Figure 10: Screenshot of Stereotype and Appropriateness questions in annotation interface.
Appendix E Feature Abbreviations
In analyses of morphosyntactic features, we consider the features listed in Masis et al. (2022). Abbreviations used for each feature are included in Table 8.
Abbrev. Feature
ZP Zero Possessive
ZC Zero Copula
DT Double Tense
HB Habitual Be
RD Resultant/Completive Done
FINNA finna
COME come
DM Double Modal
MN Multiple Negation
NAI Negative Auxiliary Inversion
NINC Non-Inverted Negative Concord
AI ain’t
3S Zero 3rd-Person Singular Present -s
IW is/was Generalization
ZPL Zero Plural -s
DO Double Object
WH Wh- Question
Table 8: Feature abbreviations and names covered by the CGEdit model Masis et al. (2022). Examples are identified from pretraining corpora using model predictions.
Appendix F Hip Hop n-gram Analysis
To estimate the inclusion of hip hop lyrics in C4, we use infinigram to efficiently count overlapping n-grams Liu et al. (2024). We first strip all capitalization and punctuation from both pretraining corpus documents (C4) and hip hop lyrics [^9] https://www.kaggle.com/datasets/nikhilnayak123/5-million-song-lyrics-dataset to account for minor differences in the presentation of the same lyrics from different sources. We then build a suffix array of the cleaned hip hop lyrics corpus using the Llama tokenizer, 16 cpus, 512MB of RAM, 1 shard. The total build time of the suffix array was approximately 10 minutes and the resulting index is approximately 6GB.
Throughout documents in the corpus, hip hop lyrics may be quoted by an author as part of a larger document. To account for this, we follow Brown et al. (2020) and consider n-gram token lengths between 8 and 13 (inclusive) to search for overlap. We exhaustively search for all occurrences of C4 n-grams in the hip hop lyrics corpus, considering any occurrence of a documents’ n-gram to be a positive label for the document as a whole. We note that this approach is not able to capture variation in orthography (e.g., spelling "going" as "goin") and relies on the comprehensiveness of the hip hop lyrics corpus; as such, it likely does not capture all hip hop lyric occurrences and underestimates actual overlap.
Appendix G Additional Hip Hop Analysis
Figure 11: Plot of vocabulary overlap between AAL-labeled C4 and Rap Lyrics corpus, with 10,000 random samples. AAL Posterior Label represents the AAL probability output by the demographic alignment classifier Blodgett et al. (2016). Lighter shades indicate a higher frequency of documents.
To complement the hip hop lyrics n-gram overlap analysis, we additionally examine the nearest neighbors of sentences in the C4 subset with the most popular hip hop songs from a separate corpus of lyrics. [^10] https://www.kaggle.com/datasets/jamiewelsh2/rap-lyrics Each song and AAL document extracted from C4 is represented by a bag-of-words (BOW) vector. We then find the nearest neighbor in the hip hop lyrics corpus by extent of unique token overlap. Figure 11 depicts the overlap in vocabulary between rap lyrics and 10,000 randomly sampled AAL texts extracted from C4. Overall, many documents share a substantial proportion of tokens with hip hop lyrics included in the lyrics corpus. In particular, 3.20% of sampled AAL texts exceed a threshold of 100 unique terms shared with hip hop lyrics.
Appendix H Non-AAL Annotations
Figure 12: Proportion of texts in the filtered (C4.en) and unfiltered (C4.en.noBlocklist - C4.en) labeled as reflecting AAL, WME, a different language variety, or no particular linguistic community.
Given the commonalities between AAL and other varieties of English (e.g., Southern White English, Oetting 2015), we then sought to investigate the texts that were labeled as AAL by the demographic-alignment classifier but not by annotators. Figure 12 presents the proportion of AAL, WME, a different variety of English, and other texts (e.g., website keywords for search engines) in the filtered and unfiltered subsets of human-annotated texts. Similar to other analyses, AAL is more represented in the unfiltered proportion of texts in C4 identified as AAL by the demographic alignment classifier. Interestingly, annotators noted a significant proportion of texts in both subsets as reflecting other varieties.
Appendix I Automated Filters
I.1 Filter Details
Descriptions of each of the 16 data filters evaluated are included in Table 9. Language filters are typically neural models used to filter datasets to a particular language (e.g., English in the case of C4 and others). Classifier-based filters train small classification models (typically using fastText Joulin et al. (2017)) or n-gram models Heafield (2011) to model text quality either through prediction of similarity to designated high-quality sources, prediction of toxicity, or n-gram model perplexity. Finally, LLM-Based filters leverage LLMs to directly predict aspects of data quality or annotate data to train a distilled model to evaluate quality.
Filter Example Use Description
fastText-langid Lui and Baldwin (2012) RefinedWeb fastText classifier trained to detect English text
fastText OH-2.5 + ELI5 Li et al. (2024) DataComp-LM* fastText model trained on the OpenHermes and ELI5 datasets
OpenWebText2 vs. RefinedWeb Li et al. (2024) DataComp-LM* fastText classifier trained with OpenWebText2 as the positive class and RefinedWeb as the negative class
Wikipedia vs. RefinedWeb Li et al. (2024) DataComp-LM* fastText classifier trained with Wikipedia as the positive class and RefinedWeb as the negative class
Multi-source vs. RefinedWeb Li et al. (2024) DataComp-LM* fastText classifier trained with Wikipedia, RedPajama, Books, and OpenWebText2 as the positive class and RefinedWeb as the negative class
Perplexity Filtering Li et al. (2024) DataComp-LM* Wikipedia-trained ngram model
fastText Toxicity Soldaini et al. (2024) Dolma fastText classifier trained on the Jigsaw toxicity dataset
PALM Quality Filter Chowdhery et al. (2022) PALM fastText classifier trained to distinguish Wikipedia, OpenWebText, and RedPajama-v1 Books as the positive class from Common Crawl.
Wikipedia-like Classifier Computer (2023) RedPajama-v1 fastText classifier trained to distinguish Wikipedia-like text from others
CC Quality Filter Gao et al. (2020) The Pile fastText classifier trained to distinguish high-quality data sources and Common Crawl in general
FineWeb-Edu Penedo et al. (2024) FineWeb-Edu Snowflake-arctic-embed-based classifier trained on Llama-3-70b-Instruct annotations of educational content
AskLLM Sachdeva et al. (2024) DataComp-LM* Prompts an LLM to measure pretraining data quality
QURating Wettig et al. (2024) QURating Prompts an LLM with 4 criteria to judge pretraining data quality
Table 9: List of the 16 evaluated pretraining corpora filters, including an example dataset or study employing each filter and a brief description. * denotes that the filter was evaluated in the development of the cited dataset, but not used in a resulting corpora. indicates filters that were replicated; all other filters are taken from their respective cited sources.
I.2 Replicated Filters
Where possible, we use filters provided by the papers that use or evaluate them. For two filters, the PALM quality filter and CC quality filter, we replicate them following available descriptions. For both filters, we train supervised fastText classifiers Joulin et al. (2017) with the default hyperparameters.
PALM Quality Filter. The PALM quality filter Chowdhery et al. (2022) uses Wikipedia samples, OpenWebText, and RedPajama-v1 books as the positive class and Common Crawl as the negative class. We randomly sample Because the books subset of RedPajama-v1 is unavailable, we instead use the Project Gutenberg books corpus.  [^11] https://huggingface.co/datasets/manu/projectgutenberg We use publicly available corpora for Wikipedia  [^12] https://huggingface.co/datasets/wikimedia/wikipedia and OpenWebText  [^13] https://huggingface.co/datasets/Skylion007/openwebtext Gokaslan et al. (2019). We sample 600 CCNet snapshots weighted among head (10%), middle (20%), and tail (70%) partitions made available through and following RedPajama-v2 Weber et al. (2024). We sample 250,000 positive and negative references, with positive references split evenly among the three corpora.
Pile-CC. The Pile-CC quality filter Gao et al. (2020) uses OpenWebText Gokaslan et al. (2019) as the positive class and Common Crawl as the negative class. Similarly to the PALM filter, we simply sample 250,000 OpenWebText and CCNet samples to form the training data.
Appendix J AAL Sources in Pretraining Corpora
Figure 13: Distribution of AAL source URL IAB categories.
We analyze the sources of AAL texts by examining the associated URL categories according to the IAB taxonomy.  [^14] https://www.iab.com/guidelines/content-taxonomy/ Using the URL’s listed in the C4 corpus, we cross reference the network location of each with a dataset of 99,015 popular website IAB classifications. [^15] https://www.kaggle.com/datasets/bpmtips/websiteiabcategorization A small proportion of texts ( 5%) are associated with URLs that are included in the dataset, suggesting that many AAL texts are not sourced from highly-visited sites. Among texts that could be matched to URL classifications, Figure 13 shows the distribution of IAB categories. A majority of texts come from Entertainment domains, containing content such as hip hop lyrics, while large proportions are also drawn from Technology & Computing domains such as social media and other forums.